find the inverse of 6x-1/2y+4=0
Answer by stanbon(75887)

find the inverse of 6x-1/2y+4=0
---------
1st: interchange x and y:
6y -(1/2)x + 4 = 0
--------
2nd: solve for "y":
6y = (1/2)x -4
---
y = (1/12)x - (2/3)
=========================
Cheers,
Stan H.